# GenVDM: Generating Vector Displacement Maps From a Single Image

### v1

@@ Abstract
We introduce the first method for generating Vector Displacement Maps (VDMs): parameterized, detailed geometric stamps commonly used in 3D modeling. Given a single input image, our method first generates multi-view normal maps and then reconstructs a VDM from the normals via a novel reconstruction pipeline. We also propose an efficient algorithm for extracting VDMs from 3D objects, and present the first academic VDM dataset. Compared to existing 3D generative models focusing on complete shapes, we focus on generating parts that can be seamlessly attached to shape surfaces. The method gives artists rich control over adding geometric details to a 3D shape. Experiments demonstrate that our approach outperforms existing baselines. Generating VDMs offers additional benefits, such as using 2D image editing to customize and refine 3D details.
@@ 2 Related work
Another line of work [74, 40] utilizes image diffusion models for novel view synthesis conditioned on an input image and a relative camera pose. Such models produce images of the object from different views, therefore the 3D object can be reconstructed by SDS-based optimization [40, 54] or a feed-forward reconstruction network [39]. These methods inspired a series of subsequent work that finetunes pretrained image diffusion models to directly generate 3D-consistent multi-view images of the target output shape given a single-view image, where the output shape can be reconstructed from generated multi-view images via optimizing a neural field or mesh [41, 60, 42], a 3D diffusion reconstruction network [38], or a feed-forward large reconstruction model powered by Transformers [27, 36, 71, 67, 89, 77, 75, 79, 73, 91]. Most recently, image diffusion models have been replaced by video diffusion models to achieve better 3D consistency of the generated views [24, 68].
@@ 3.2 VDM Reconstruction
Reconstructing 3D shapes from multi-view images has been well studies in the text/image-to-3D literature. Most recent methods adopt a feed-forward large reconstruction model (LRM) to directly generate a 3D shape from multiple input images of different viewpoints [27, 36, 71, 67, 89, 75]. Therefore, a straightforward way for reconstructing VDMs is to train a similar LRM to take the normal maps as input and directly regress a VDM image. However, given limited VDM training shapes, our LRM trained on a small dataset is unlikely to generalize as well as other LRM models trained on larger datasets, therefore leading to suboptimal results.
@@
$\begin{split}\operatorname*{argmin}_{\theta}\;\;\mathbb{E}_{P,Q}\;&\frac{1}{|P|}\sum_{p\in P}\min_{q\in Q}\|\phi_{\theta}(p)-q\|_{2}^{2}+\\ &\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in P}\|\phi_{\theta}(p)-q\|_{2}^{2}+\\ &\frac{1}{|\partial P|}\sum_{p\in\partial P}\|\phi_{\theta}(p)-\text{proj}(p)\|_{2}^{2},\end{split}$ (1)
@@ 3.3 Data Preparation
$\operatorname*{argmin}_{A^{\prime}}\;\;\mathbb{E}_{(p,q)\in E}\;\|(p^{\prime}-q^{\prime})-(p-q)\|_{2}^{2}.$ (2)
The minimization objective is to ensure that the gradients on the mesh are preserved as much as possible after deformation, while the target coplanar boundary points $B^{\prime}$ are also strictly followed.
@@ 4.1 Vector Displacement Map Generation
Evaluation Dataset and Metrics.
As there is no existing benchmark dataset for VDMs, we collected a dataset of 50 RGB images from the Internet and a text-to-image model [5] for evaluation. All images depict common VDM categories used by artists such as facial elements and decorations. For quantitative evaluation, we measure CLIP similarity [55] and 3D-FID score [72] between the input image and the rendered images of the generated shapes from different views, denoted as CLIPImg and 3D-FID, respectively. For CLIP, we additionally assess semantic alignment by measuring CLIP similarity between rendered images and texts describing the categories of the input images, denoted as CLIPText. We use public implementation of CLIP [62] and 3D-FID [57] for computing the metrics. Please see Supplementary Material for more details.
Figure 7: Qualitative results of ablation study.
4.2 Ablation Study
@@
Table 1: Quantitative comparison with baseline methods. Scalar DM stands for scalar displacement map produced from DepthAnything [84].
@@ Appendix C Details on VDM Reconstruction
where $P$ are sets of sampled points from $[0,1]^{2}$ and $(p)$ maps p to the corresponding 3D point in a pre-defined 3D square plane. We then use the optimization objective proposed in 3.2 for subsequent optimization. For mesh optimization comparison method, we set the laplacian regularization loss ratio to 1e-4 compared with chamfer reconstruction loss. We optimize for 1200 iterations and no remeshing is done during the optimization process. For topological fixing and tutte embedding comparison method, we use the built-in parametrization algorithm in Blender[1] ”Unwrap” function in UV editing after fixing boundary vertices.
@@ Appendix D Details on Training
Appendix E More Result

### v2

@@ Abstract
We introduce the first method for generating Vector Displacement Maps (VDMs): parameterized, detailed geometric stamps commonly used in 3D modeling. Given a single input image, our method first generates multi-view normal maps and then reconstructs a VDM from the normals via a novel reconstruction pipeline. We also propose an efficient algorithm for extracting VDMs from 3D objects, and present the first academic VDM dataset. Compared to existing 3D generative models focusing on complete shapes, we focus on generating parts that can be seamlessly attached to shape surfaces. The method gives artists rich control over adding geometric details to a 3D shape. Experiments demonstrate that our approach outperforms existing baselines. Generating VDMs offers additional benefits, such as using 2D image editing to customize and refine 3D details. Code and data are available at https://yyuezhi.github.io/GenVDM/.
@@ 2 Related work
Another line of work [74, 40] utilizes image diffusion models for novel view synthesis conditioned on an input image and a relative camera pose. Such models produce images of the object from different views, therefore 3D object can be reconstructed by SDS-based optimization [54] or a feed-forward reconstruction network [39]. These methods inspired a series of subsequent work that finetunes pretrained image diffusion models to directly generate 3D-consistent multi-view images of target output shape given a single-view image, where output shape can be reconstructed from generated multi-view images via optimizing a neural field or mesh [41, 60, 42], a 3D diffusion reconstruction network [38], or a feed-forward large reconstruction model powered by Transformers [27, 36, 71, 67, 89, 77, 79, 73, 91]. Most recently, image diffusion models have been replaced by video diffusion models to achieve better 3D consistency of generated views [24, 68].
@@ 3.2 VDM Reconstruction
Reconstructing 3D shapes from multi-view images has been well studied in the text/image-to-3D literature. Most recent methods adopt a feed-forward large reconstruction model (LRM) to directly generate a 3D shape from multiple input images of different viewpoints [27, 36, 71, 67, 89, 75]. Therefore, a straightforward way for reconstructing VDMs is to train a similar LRM to take the normal maps as input and directly regress a VDM image. However, given limited VDM training shapes,
our LRM trained on a small dataset is unlikely to generalize as well as other LRM models trained on larger datasets, therefore leading to suboptimal results.
@@
$\begin{split}\operatorname*{argmin}_{\theta}\;\;\mathbb{E}_{P,Q}\;&\frac{1}{|P|}\sum_{p\in P}\min_{q\in Q}\|\phi_{\theta}(p)-q\|_{2}^{2}+\frac{1}{|Q|}\sum_{q\in Q}\min_{p\in P}\|\phi_{\theta}(p)-q\|_{2}^{2}+\\ &\frac{1}{|\partial P|}\sum_{p\in\partial P}\|\phi_{\theta}(p)-\text{proj}(p)\|_{2}^{2},\end{split}$
@@ 3.3 Data Preparation
$\operatorname*{argmin}_{A^{\prime}}\;\;\mathbb{E}_{(p,q)\in E}\;\|(p^{\prime}-q^{\prime})-(p-q)\|_{2}^{2}.$
The minimization objective is to ensure that the gradients on the mesh are preserved after deformation, while the target coplanar boundary points $B^{\prime}$ are also strictly followed.
@@ 4.1 Vector Displacement Map Generation
Table 1: Quantitative comparison with baselines. Scalar DM is scalar displacement map produced from DepthAnything [84].
@@
Evaluation Dataset and Metrics.
As there is no existing benchmark dataset for VDMs, we collected a dataset of 50 RGB images from the Internet and a text-to-image model [5] for evaluation. All images depict common VDM categories used by artists such as facial elements and decorations. For quantitative evaluation, we measure CLIP similarity [55] and 3D-FID score [72] between the input image and the rendered images of the generated shapes from different views, denoted as CLIPImg and 3D-FID, respectively. For CLIP, we additionally assess semantic alignment by measuring CLIP similarity between rendered images and texts describing the categories of the input images, denoted as CLIPText. We use public implementation of CLIP [62] and 3D-FID [57] for computing the metrics.
Figure 7: Qualitative results of ablation study.
4.2 Ablation Study
@@ 5 Conclusion, Limitation, and Future Work
Acknowledgment. Yuezhi Yang and Qixing Huang were supported by NSF IIS-2047677, NSF IIS-2413161, and Gifts from Adobe and Google.
@@ Appendix C Details on VDM Reconstruction
where $P$ are sets of sampled points from $[0,1]^{2}$ and $(p)$ maps p to the corresponding 3D point in a pre-defined 3D square plane. We then use the optimization objective proposed in 3.2 for subsequent optimization. For mesh optimization comparison method, we set the laplacian regularization loss ratio to 1e-4 compared with chamfer reconstruction loss. We optimize for 1200 iterations and no remeshing is done during the optimization process. For topological fixing and tutte embedding comparison method, we use the built-in parametrization algorithm in Blender[1] "Unwrap" function in UV editing after fixing boundary vertices.
@@ Appendix D Details on Training
Appendix E Additional Evaluation
We added 50 RGB images from the Internet and a text-to-image model [5] into our testing dataset of 50 images used in the paper. Quantitative results on the new 100-image dataset are shown in Table 3.
Method CLIPImg↑ CLIPText↑ 3D-FID↓
Wonder3D [42] 0.8143 0.2504 208.7
Magic123 [54] 0.8241 0.2488 218.5
LRM [27] 0.8122 0.2472 246.3
Scalar DM 0.8201 0.2529 223.4
Ours 0.8460 0.2639 197.3
Table 3: Additional quantitative comparison with baselines.
Appendix F More Result